OpenStudy (anonymous):

**Calc 2 Help** let u∠ = 2i∠ + 3j∠ , and v∠ = 8i∠ - 13j∠ , and w∠ = -i∠ + j∠ . Find real numbers a and b so that u∠ =av∠ + bw∠ .

OpenStudy (anonymous):

The "∠ " is mean to be over the preceeding character.

OpenStudy (amistre64):

its best just to write it out in component form ... u=<2,3> v = <8,-13> w=<-1,1>

OpenStudy (amistre64):

the rest of it just amounts to distractors

OpenStudy (amistre64):

this is also a matrix solution ....

OpenStudy (anonymous):

should I do something like this.. <2,3> = a<8,-13> + b<-1,1>?

OpenStudy (amistre64):

A x = v | 8 -1| |a| = |2| |-13 1 | |b| |3|

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

8a -b = 2 -13a +b = 3 which defines a system of equation

OpenStudy (anonymous):

@m.f.h sorry I wasn't sure how to type the notation.

OpenStudy (anonymous):

would it be <2,3> = <8a,-13a> + <-b,b>?

OpenStudy (anonymous):

i know ∠ mean the angle you can find this notation in calculator

OpenStudy (amistre64):

yes, it would amount to that

OpenStudy (amistre64):

add up the parts and compare them to the results your looking to get

OpenStudy (amistre64):

<8a,-13a> + <-b, b > ------------- <8a-b,-13a+b> = <2 , 3 >

OpenStudy (amistre64):

8a -b = 2 -13a +b = 3 defines a system of equation which shouldve been lerant back in algebra classes

OpenStudy (anonymous):

I solved for b 8a -2 =b, b = 3 +13a 8a-2 = 3 +13a -5 = 5a a = -1

OpenStudy (anonymous):

is that not how I should do it?

OpenStudy (amistre64):

as a dbl chk when your ready rref{{8,-1,2},{-13,1,3}} http://www.wolframalpha.com/input/?i=rref%7B%7B8%2C-1%2C2%7D%2C%7B-13%2C1%2C3%7D%7D a does = -1, good job so far

OpenStudy (amistre64):

if your comfortable with substitution, then by all means use it; its fine to go that route

OpenStudy (anonymous):

ok b = -10